A boat is 9km north then 15 km west from the starting point. How far is the boat from its starting point?
Answer by mananth(16946)

North and west are at right angles
so pythagoras theorem will apply
the distance from starting point to the final point acts as the hypotenuse.
so apply pythagoras theorem
d^2= 15^2+9^2
d^2= 225+81
d^2= 306
d= sqrt(306)
d= 17.5 km
the distance from the starting point